Consider the following set of data.

(21, 5), (33, 51), (59, 32), (83, 21), (113, 58), (119, 8)

(a) Calculate the covariance of the set of data. (Give your answer correct to two decimal places.)


(b) Calculate the standard deviation of the six x-values and the standard deviation of the six y-values. (Give your answers correct to three decimal places.)

(c) Calculate r, the coefficient of linear correlation, for the data in part (a). (Give your answer correct to two decimal places.)
Answer #1

Solution :

x y (x-xbar) (y-ybar) (x-xbar)(y-ybar) (x-xbar)^2 (y-ybar)^2
21 5 -50.3333 -24.1667 1216.389 2533.444 584.0278
33 51 -38.3333 21.83333 -836.944 1469.444 476.6944
59 32 -12.3333 2.833333 -34.9444 152.1111 8.027778
83 21 11.66667 -8.16667 -95.2778 136.1111 66.69444
113 58 41.66667 28.83333 1201.389 1736.111 831.3611
119 8 47.66667 -21.1667 -1008.94 2272.111 448.0278
Sum 428 175 0 0 441.6667 8299.333 2414.833
Mean xbar = 71.33333 ybar = 29.16667

a)

Covariance = \frac{\sum (x-\bar{x})(y-\bar{y})}{n-1} = \frac{441.6667}{6-1} = \frac{441.6667}{5} = \mathbf{88.33}

b)

Standard \;deviation\; of\; x\;\; S_{x} =\sqrt \frac{\sum (x-\bar{x})^2}{n-1} = \sqrt \frac{8299.333}{6-1} = \sqrt\frac{8299.333}{5}

S_{x} = \mathbf{40.741}

Standard \;deviation\; of\; y\;\; S_{y} =\sqrt \frac{\sum (y-\bar{y})^2}{n-1} = \sqrt \frac{2414.833}{6-1} = \sqrt\frac{2414.833}{5}

S_{y} = \mathbf{21.977}

c)

r = \frac{Covariance(x,y)}{Standard \;deviation \;of\; x * Standard\; deviation\; of\; y} = \frac{Cov(x,y)}{S_{x}*S_{y}}

r = \frac{88.33}{40.741 * 21.977} = \frac{88.33}{895.365} = \mathbf{0.10}